### v1

# ST-FlowNet: An Efficient Spiking Neural Network for Event-Based Optical Flow Estimation

Abstract
Spiking Neural Networks (SNNs) have emerged as a promising tool for event-based optical flow estimation tasks due to their ability to leverage spatio-temporal information and low-power capabilities. However, the performance of SNN models is often constrained, limiting their application in real-world scenarios. In this work, we address this gap by proposing a novel neural network architecture, ST-FlowNet, specifically tailored for optical flow estimation from event-based data. The ST-FlowNet architecture integrates ConvGRU modules to facilitate cross-modal feature augmentation and temporal alignment of the predicted optical flow, improving the network’s ability to capture complex motion dynamics. Additionally, to overcome the challenges associated with training SNNs, we introduce a novel approach to derive SNN models from pre-trained artificial neural networks (ANNs) through ANN-to-SNN conversion or our proposed BISNN method. Notably, the BISNN method alleviates the complexities involved in biological parameter selection, further enhancing the robustness of SNNs in optical flow estimation tasks. Extensive evaluations on three benchmark event-based datasets demonstrate that the SNN-based ST-FlowNet model outperforms state-of-the-art methods, delivering superior performance in accurate optical flow estimation across a diverse range of dynamic visual scenes. Furthermore, the inherent energy efficiency of SNN models is highlighted, establishing a compelling advantage for their practical deployment. Overall, our work presents a novel framework for optical flow estimation using SNNs and event-based data, contributing to the advancement of neuromorphic vision applications.
keywords:
Spiking neural networks , Optical flow estimation , Event-based images , Training methods.
1 Introduction
As the third generation of neural networks (Maass, 1997), spiking neural networks (SNNs) have garnered increasing attention in recent years (Zhang et al., 2022; Luo et al., 2023; Zheng et al., 2021). Unlike conventional artificial neural networks (ANNs), SNNs employ bio-inspired neurons and discrete spike trains to mimic the complex spatiotemporal dynamics of the human brain (Zheng et al., 2024; Sun et al., 2024). These characteristics enable SNNs to achieve competitive performance across a wide range of neuromorphic intelligent tasks (Ussa et al., 2024; Hagenaars et al., 2021; Yu et al., 2023a). Furthermore, the event-driven information coding and transmission mechanisms ensure lower power consumption, thus enhancing the feasibility of SNNs for hardware implementation (Han and Roy, 2020; Yao et al., 2023).
Optical flow estimation is a fundamental topic in the field of computer vision and has extensive applications (Fleet and Weiss, 2006; Ilg et al., 2017; Vandal and Nemani, 2023), particularly in motion-related intelligent tasks such as object segmentation (Zitnick et al., 2005), video enhancement (Fan et al., 2021; Wang et al., 2020), and object tracking (Kale et al., 2015; Ussa et al., 2024). However, prevalent research has predominantly concentrated on optical flow estimation from frame-based images captured by conventional cameras, resulting in significant performance degradation in challenging scenarios characterized by high-speed motion or unfavorable lighting conditions (Otte and Nagel, 1994; Zhai et al., 2021). Event-based neuromorphic cameras, capable of asynchronously recording changes in light intensity within a high dynamic range of illumination (Han et al., 2020), respond to these challenging scenarios effectively and exhibit potential for minimizing energy consumption, thereby presenting an attractive prospect for deployment on edge devices (Yu et al., 2023a). Furthermore, drawing inspiration from inherent imaging mechanisms, event-based images can circumvent errors introduced by assumptions related to the conservation of pixel intensity in optical flow estimation (Gaur, 2022; Barron et al., 1994). Thus, the advantages of event-based optical flow estimation are significant when compared with conventional approaches (Benosman et al., 2012; Zhu et al., 2018b; Mueggler et al., 2017; Scheerlinck et al., 2020).
Figure 1: The framework of our proposed optical flow estimation method is illustrated. The ST-FlowNet (both ANN and SNN) models utilize event-based images as input data. Following training on the ANN model, an SNN ST-FlowNet model is derived through the A2S conversion or BISNN method. Optical flow prediction is achievable using both ANN and SNN models. Additionally, for reference, the corresponding frame-based images are presented in the left black box. Conventional frame-based images exhibit abundant spatial texture information indiscriminately, while event-based images emphasize motion-related objects by leveraging spatio-temporal cues simultaneously.
Owing to the structural specificity inherent in the event modality, current approaches typically involve the reconstruction of event data into frame-based images for the estimation tasks handled by conventional ANN models (Zhu et al., 2018a; Dosovitskiy et al., 2015; Ilg et al., 2017). Despite some notable progress, these approaches often neglect potential errors inherent in the reconstruction process. Representatively, the resulting reconstructed images frequently exhibit pronounced motion blur in spatial visual cues, particularly as the sampling time window lengthens (Hagenaars et al., 2021; Stoffregen et al., 2020; Tian and Andrade-Cetto, 2022). Concurrently, the valuable dynamic features embedded in the temporal domain remain underutilized. In contrast, an SNN model, comprising spiking neurons as fundamental units, accepts spiking events as input and uses firing spikes as a medium for information propagation and presentation (Ghosh-Dastidar and Adeli, 2009; Tavanaei et al., 2019; Cai et al., 2024). The SNN is an essential tool for event-based feature extraction (Sun et al., 2024; Yu et al., 2022). Consequently, a natural solution to address the aforementioned challenges involves developing an SNN model specifically tailored for event-based optical flow estimation. However, training SNNs using the standard backpropagation (BP) method is hindered by the non-differentiable nature of the spikes (Wu et al., 2018), resulting in constrained performance for real-world tasks.
To fully exploit the potential of event data in optical flow estimation, it is imperative to effectively address two key challenges: (1) developing a novel architecture capable of simultaneous spatio-temporal feature extraction based on event data; and (2) proposing a novel model training strategy for achieving superior optical flow estimation performance. Thus far, limited research has delved into these pertinent issues. Therefore, this work introduces a novel method to effectively tackle these challenges (shown in Fig. 1). We present our ST-FlowNet architecture tailored specifically to estimate event-based optical flow. By incorporating the ConvGRU layers (Ballas et al., 2016), ST-FlowNet achieves cross-scale fusion of dynamic optical flow features from event data in the temporal dimension. In contrast to prior models inspired by pyramidal architectures (Dosovitskiy et al., 2015; Ilg et al., 2017), ST-FlowNet employs a more streamlined decoder structure, enabling direct decoding of latent information across the entire multi-scale feature space. Following training on the ANN model, the SNN model for optical flow estimation can be derived through two strategies: the ANN-to-SNN conversion method (A2S) or our proposed bio-information-fused training strategy (BISNN). Notably, our BISNN approach achieves parameter-free model conversion while preserving the performance of optical flow estimation.
Our key contributions in this paper are summarized as follows:
  • 1.
    We propose ST-FlowNet, a novel architecture designed for efficient optical flow estimation by leveraging spatio-temporal features in event-based data.
  • 2.
    We present the first framework for constructing efficient SNN models for optical flow estimation, utilizing an A2S conversion approach.
  • 3.
    We introduce a novel parameter-free SNN training strategy, which further mitigates training challenges while enhancing overall training efficiency.
  • 4.
    Our experimental results demonstrate that ST-FlowNet attains superior performance when compared with other state-of-the-art models on challenging benchmark datasets such as MVSEC, ECD, and HQF.
2 Related Work
Optical flow estimation is a fundamental task in the field of computer vision (Fleet and Weiss, 2006; Ilg et al., 2017), garnering significant research attention. In this section, we initially delve into the evolution of event-based optical flow estimation. Subsequently, our focus shifts to the models built by SNNs, which demonstrate notable proficiency in extracting spatio-temporal features from event-based images.
2.1 Event-based Optical Flow Estimation
Considering the distinctive attributes of event images in contrast to the RGB modality, early research focused on innovating paradigms for event-based optical flow computation, but only achieved limited success in several simple scenarios (Benosman et al., 2012, 2013). The introduction of large-scale benchmark datasets has facilitated applying deep learning models to event-based optical flow estimation tasks (Zhu et al., 2018b; Mueggler et al., 2017; Scheerlinck et al., 2020), significantly enhancing the efficacy of optical flow estimation and reducing the challenges associated with model design (Scheerlinck et al., 2020; Tian and Andrade-Cetto, 2022; Zhu et al., 2018a). Previous research attempted to use encoder-decoder networks to decouple spatio-temporal features embedded in event data across multiple resolutions (Zhu et al., 2018a). A series of variants emerged through the refinement of network structures and adjustments to loss functions, achieving reliable and accurate optical flow estimation facilitated by self-supervised and end-to-end learning methods (Scheerlinck et al., 2020; Tian and Andrade-Cetto, 2022). However, considering the high temporal resolution characteristics inherent in event data, models founded on general convolution structures struggle to comprehensively use the temporal features of the event data, thereby limiting the effectiveness of event-based optical flow estimation (Rueckauer and Delbruck, 2016).
2.2 SNN Models for Optical Flow Estimation
Considering the inherent advantages of SNNs in spatio-temporal feature extraction (Ghosh-Dastidar and Adeli, 2009; Tavanaei et al., 2019), our work focuses on models using spiking neurons which are naturally adept at capturing spatio-temporal visual cues embedded in event data for precise optical flow estimation. Recently, numerous models based on spiking neurons have been proposed for event-based optical flow estimation (Diehl and Cook, 2015; Paredes-Vallés et al., 2019; Orchard et al., 2013; Haessig et al., 2018; Hagenaars et al., 2021; Zhang et al., 2023). Similar to traditional ANN models, shallow convolutional SNN models trained using spike-time-dependent-plasticity learning methods (Diehl and Cook, 2015) have been introduced, demonstrating promising performance for relatively simple tasks (Paredes-Vallés et al., 2019; Orchard et al., 2013; Haessig et al., 2018). To enhance SNN model performance in complex real-world scenarios, a logical approach involves constructing hybrid models that leverage the strengths of both ANNs and SNNs (Diehl and Cook, 2015). Additionally, drawing inspiration from the BP-style direct training method for SNN models (Wu et al., 2018), a series of fully spiking models have also attained state-of-the-art performance on event-based data when compared with advanced ANN models (Hagenaars et al., 2021; Zhang et al., 2023). However, constrained by training approaches and network architectures, we posit that event-based optical flow estimation by SNN models can be further optimized by ANN conversion (Deng and Gu, 2021; Li et al., 2022).
3 Method
We present an overview of the proposed ST-FlowNet model in Fig. 2. Specifically, ST-FlowNet builds on the foundation of the end-to-end FlowNet model (Dosovitskiy et al., 2015; Ilg et al., 2017), characterized by encoder and decoder layers. To effectively leverage the event modality, which inherently possesses rich spatio-temporal information, we incorporate ConvGRU layers (Cho et al., 2014) for sequential feature argumentation and alignment. Additionally, we use SNN models derived from pre-trained ANN models to enhance the ST-FlowNet’s proficiency in spatio-temporal feature extraction. In this section, we elaborate on the event representation, network architecture, and training methodology used for ST-FlowNet.
Figure 2: (a) The ST-FlowNet architecture is illustrated. Following pre-processing by a ConvGRU layer, the enhanced event-based input undergoes downsampling via four encoder layers. The resulting minimal feature maps produced by encoder4 traverse two residual block layers, ensuring robust feature extraction. Through the concatenation of feature maps at various levels, decoder layers and a generator are deployed for basic optical flow prediction. Furthermore, the basic predicted optical flow is fed through a ConvGRU layer to fuse historical sequential temporal feature and generate the final predicted optical flow. (b) Schematic illustration depicting the architecture of ConvGRU. A ConvGRU unit integrates both current input and state information to produce a corresponding output. The symbol $\odot$ denotes the Hadamard product, and $\sigma$ is the activation function.
3.1 Event Representation
In contrast to conventional frame-based images, which capture light intensity at discrete exposure times, event-based images ($\left\{\left[x_{k},y_{k},t_{k},p_{k}\right]\right\}_{k=1}^{K}$) asynchronously record changes in light intensity at location $[x_{k},y_{k}]$ along with their respective polarity ($p_{k}\in\left\{p^{+},p^{-}\right\}$). Here, $K$ represents the total number of events, and $t_{k}$ denotes the timestamp of the $k$-th event. For the sake of convenience, the event inputs are aggregated into $N$ group of frame-based representations $f_{n}^{+}(\tilde{x},\tilde{y})$ and $f_{n}^{-}(\tilde{x},\tilde{y})$ ($n=1,...,N$), which are defined by the following formulas:
$f_{n}^{+}(\tilde{x},\tilde{y})$ $={\textstyle\sum_{i}}p_{i}^{+},$ (1)
$f_{n}^{-}(\tilde{x},\tilde{y})$ $={\textstyle\sum_{i}}p_{i}^{-},$ (2)
subject to $i=\frac{K}{N}(n-1)+1,...,\frac{K}{N}n$. The inputs to the models are presented as:
[…]
Concretely, we concatenate N groups of event-based images into a $2N$-channel data representation (Eq. 3), which serves as input to the ANN model, ensuring the complete information of the inputs is captured. For the SNN model, the value of N determines the temporal resolution of the raw data, and each pair of images is sequentially processed by the model (Eq. 4).
3.2 Network Architecture
As illustrated in Fig. 2, the ST-FlowNet model is structured in a semi-pyramidal encoder-decoder architecture (Dosovitskiy et al., 2015; Ilg et al., 2017). Diverging from prior approaches, we incorporate ConvGRU layers (Cho et al., 2014) to enhance the model’s sequential feature representation. The ConvGRU layer retains its original architecture (shown in Fig. 2(b)). However, the hidden state, which is used to modify information across different time steps, has been redesigned.
Specifically, to process the input data, a ConvGRU layer is employed to combined the present input ${I}_{t}$ with the inherent information $O_{t-1}$. The augmented input data $\tilde{I}_{t}$ is characterized as follows:
$\tilde{I}_{t}=(1-Z_{t})\odot O_{t-1}+Z_{t}\odot H_{t}.$ (5)
$R_{t}$, $Z_{t}$ and $H_{t}$ represent the reset gate, update gate and candidate hidden state respectively, and are calculated as follows:
$R_{t}$ $=\mathrm{\sigma_{1}}(\mathrm{\xi_{R}}(O_{t-1},I_{t})),$ (6)
$Z_{t}$ $=\mathrm{\sigma_{1}}(\mathrm{\xi_{Z}}(O_{t-1},I_{t})),$ (7)
$H_{t}$ $=\mathrm{\sigma_{2}}(\mathrm{\xi_{I}}(R_{t}\odot O_{t-1},I_{t})).$ (8)
Here, $\mathrm{\xi_{R}}$, $\mathrm{\xi_{Z}}$ and $\mathrm{\xi_{I}}$ represent convolutional operations, $\mathrm{\sigma_{1}}$ and $\mathrm{\sigma_{2}}$ are the sigmoid and tanh activation functions. Intuitively, the candidate hidden state $H_{t}$ balances and retains pertinent information from the previous optical flow and current input (8). The reset gate $R_{t}$ governs the degree to which historical information is removed in $H_{t}$. Ultimately, the augmented input $\tilde{I}_{t}$ undergoes an update via a weighted summation of $O_{t-1}$ and $H_{t}$ (5). $O_{-1}$ is established as 00 at the initial time step $t=0$.
For every augmented input $\tilde{I_{t}}$, ST-FlowNet uses a convolutional architecture as Encoders to derive feature maps across multiple hierarchical levels, spanning from low-level to high-level representations (denoted as $\{F_{t}^{l}\}_{l=1}^{4}$). The dimensions of the feature maps undergo a reduction by half in each subsequent layer, while the channel count experiences a twofold increment. Furthermore, the resultant highest-level representation $F_{t}^{4}$ undergoes processing via two residual Block, aimed at conducting a more in-depth exploration of the underlying deep features (denoted as $\tilde{F_{t}^{4}}$).
To maximize the exploitation of optical flow information across various spatial scales, our architecture concatenates all feature maps that have undergone downsampling to attain a uniform size. Subsequently, this concatenated representation is upsampled to yield an input for subsequent decoder stages. This distinctive procedure is formally written as follows:
$F_{t}=\mathcal{U}_{2}(\mathrm{concate}(\mathcal{D}_{8}(F_{t}^{1}),\mathcal{D}_{4}(F_{t}^{2}),\mathcal{D}_{2}(F_{t}^{3}),\tilde{F_{t}^{4}})),$ (9)
where $\mathcal{D}$ and $\mathcal{U}$ represent the down and upsampling operations, respectively, with a subscript denoting the specific sampling factor. In our approach, we implement the downsampling process using convolutional layers with corresponding strides, while the upsampling process is achieved through bilinear interpolation.
After the feature extraction stage, tandem Decoder modules, each consisting of a convolutional layer with uniform input and output dimensions, decode the cross-scale feature representation into a predicted optical flow. The decoders are tasked with interpreting the flow information embedded in the feature space and generating a corresponding flow map. Finally, a Generator module, comprising a single convolutional layer with two filters, produces the basic predicted flow. In our method, the generator acts as a refinement stage, further enhancing the optical flow prediction generated by the decoders.
To further improve the optical flow prediction, we incorporate a ConvGRU2 layer (Cho et al., 2014) following the Generator module to output the final optical flow. ConvGRU2 uses the predicted flow from the previous time step as its state and the basic predicted flow from the current time step as its input. This temporal integration of optical flow information enables ST-FlowNet to capture long-range dependencies in the flow sequence. In addition, the output $O_{t}$ of ConvGRU2 is then used to augment the state information for ConvGRU1 at the next time step, providing a more comprehensive representation of the optical flow dynamics.
3.3 Model Training
In this work, we use two types of cross-model transformation methodology, A2S conversion (Bu et al., 2022; Deng and Gu, 2021; Rathi and Roy, 2020; Ding et al., 2021; Deng et al., 2020) and bio-information fusion training, to generate an optimal SNN for optical estimation while preserving the original ANN model’s accuracy.
3.3.1 ANN-to-SNN conversion
The fundamental process of the A2S conversion encompasses three key phases: first, constructing ANN and SNN models with identical architectures but employing distinct basic neurons; second, training ANN models using the standard BP method; and third, converting the trained ANN model into an SNN model, ensuring extensive retention of model accuracy.
In the ANN model, where information is represented with continuous values, the neuron is defined as follows:
$a^{l}=h(w^{l}\cdot a^{l-1}),$ (10)
where $a^{l}$ is the output in the $l$-th layer, $w^{l}$ signifies the weight between the $(l-1)$-th and $l$-th layers, and $h(\cdot)$ denotes the activation function. In the SNN model, we use the leaky integrate-and-fire model (Abbott, 1999) (LIF) as the spiking neuron for network construction. Without loss of generality, the iterative form of the membrane potential $v^{l}(t)$ of the LIF neuron is described as follows (Rathi and Roy, 2020; Han and Roy, 2020):
$v^{l}(t)=e^{-\tau}\cdot v^{l}(t-1)-s^{l}(t-1)\cdot\theta^{l}+w^{l}\cdot s^{l-1}(t)\cdot\theta^{l-1}.$ (11)
Here, $\theta^{l}$ is the spiking firing threshold, $\tau\in\left[0,+\infty\right)$ represents the membrane potential decay factor, and $s^{l}$ denotes the spike output generated by:
$s^{l}(t)=H(v^{l}(t)-\theta^{l}).$ (12)
Using the Heaviside step function $H(\cdot)$, the LIF spiking neuron emits a spike once the membrane potential surpasses the predetermined firing threshold. To circumvent notable accuracy diminution, we use the quantization clip-floor activation function (Bu et al., 2022) in the ANN model instead of the ReLU function:
$h(x)=\lambda^{l}\cdot clip(\frac{1}{L}\left\lfloor\frac{x\cdot L}{\lambda^{l}}\right\rfloor,0,1).$ (13)
[…]
The training process for the optical flow estimation network is conducted using the ANN model. Motivated by the limited availability of ground truth optical flow data, we train our optical flow estimation network through a self-supervised approach (Hagenaars et al., 2021). The comprehensive loss function encompasses two fundamental components: a contrast loss $\mathcal{L}_{\mathrm{contrast}}$ and a smoothness loss $\mathcal{L}_{\mathrm{smooth}}$. The contrast loss $\mathcal{L}_{\mathrm{contrast}}$ uses a reformulated contrast maximization proxy loss to gauge the accuracy of optical flow estimation by assessing the motion compensation performance of an image reconstructed from the predicted optical flow (Hagenaars et al., 2021; Gallego et al., 2019, pp. 12280–12289; Mitrokhin et al., 2018; Zhu et al., 2019). The smoothness loss $\mathcal{L}_{\mathrm{smooth}}$ uses Charbonneir smoothness function (Charbonnier et al., 1994; Zhu et al., 2018a, 2019) to regulate the optical flow variation between neighboring pixels. Consequently, the total loss is defined as:
$\mathcal{L}_{\mathrm{flow}}=\mathcal{L}_{\mathrm{contrast}}+\lambda\mathcal{L}_{\mathrm{smooth}},$ (14)
where the scalar $\lambda$ balances the respective weights of the contrast loss and the smoothness loss.
Once the ANN model is fully trained, the network weights $\left\{w^{l}\right\}_{l=0}^{L}$ and parameters $\left\{\lambda^{l}\right\}_{l=0}^{L}$ of the ANN are transformed into the weights and spiking fire thresholds $\left\{\theta^{l}\right\}_{l=0}^{L}$ in the SNN (Deng and Gu, 2021; Deng et al., 2020).
3.3.2 Bio-information-fused training
Although the A2S conversion demonstrates superior performance in previous work, there are still some inherent problems that are hard to avoid. The biology parameters, spiking firing threshold and membrane potential decay factor in LIF neuron, significantly influence the spatio-temporal information processing capability of SNN models (Yu et al., 2023b; Sun et al., 2024; Fang et al., 2021). In A2S conversion method, these parameters are usually determined empirically or by threshold balancing strategies, limiting the performance of converted models.
To address these challenges, we propose a bio-information-fused training strategy (BISNN). This approach incorporates two key operations: (1) Cross-model initialization: The SNN models, which share an identical architecture, are initialized with the pre-trained weights $\left\{w^{l}\right\}_{l=0}^{L}$ of the ANN models. This facilitates efficient knowledge transfer between models; and (2) Parameter-free optimization: A supervised retraining procedure utilizing the STBP (Wu et al., 2018) method is employed to optimize the SNN models, thereby circumventing the need for complex biological parameter filtering processes. According to the chain rule, the mathematical formulation of the loss function’s derivatives with respect to the learnable parameters $\left\{w^{l}\right\}_{l=0}^{L}$ can be expressed as follows:
$\frac{\partial\mathcal{L}_{\mathrm{flow}}}{\partial v^{l}(t)}=\frac{\partial\mathcal{L}_{\mathrm{flow}}}{\partial s^{l}(t)}\frac{\partial s^{l}(t)}{\partial v^{l}(t)}+\frac{\partial\mathcal{L}_{\mathrm{flow}}}{\partial s^{l}(t+1)}\frac{\partial s^{l}(t+1)}{\partial v^{l}(t)},$ (15)
$\frac{\partial\mathcal{L}_{\mathrm{flow}}}{\partial w^{l}}=\sum_{t=1}^{T}\frac{\partial\mathcal{L}_{\mathrm{flow}}}{\partial v^{l}(t)}\frac{\partial v^{l}(t)}{\partial w^{l}}.$ (16)
In this study, the approximate gradient function employed in the backpropagation process is formulated as follows (Fang et al., 2021):
$H(\cdot)=\frac{\text{arctan}[\pi(\cdot)]}{\pi}+\frac{1}{2},$ (17)
where arctan represents the inverse tangent function.
Table 1: Comparison of event-based datasets for optical flow estimation. GT denotes the ground truth datasets.
Datasets Year Resolution Train Test GT
UZH-FPV (Zhu et al., 2018b) 2018 346$\times$260 $\times$ $\times$
ECD (Mueggler et al., 2017) 2017 240$\times$180 $\times$ $\times$
MVSEC (Zhu et al., 2018b, a) 2019 346$\times$260 $\times$
HQF (Stoffregen et al., 2020) 2020 240$\times$180 $\times$ $\times$
Table 2: Comparison of state-of-the-art models across different datasets. The ST-FlowNet model trained using the ANN, A2S, and BISNN methods are denoted as ST-FlowNet1, ST-FlowNet2, and ST-FlowNet3, respectively. For the MVSEC dataset, the AEE results for each scenario are presented. For the ECD and HQF datasets, the average FWL and RSAT results across all scenarios are reported. The best results are highlighted in bold. The symbol $\downarrow$ ($\uparrow$) indicates that a smaller (larger) value is preferred.
Model OD1 IF1 IF2 IF3 ECD HQF
[…]
ANN EV-FlowNet (Hagenaars et al., 2021) 0.32 1.30 0.58 2.18 1.02 3.85 0.87 3.18 1.31 0.94 1.37 0.92
GRU-EV-FlowNet (Hagenaars et al., 2021) 0.47 1.69 0.60 2.16 1.17 3.90 0.93 3.00
GRU-FireNet (Scheerlinck et al., 2020) 0.55 2.04 0.89 3.35 1.62 5.71 1.35 4.68
ET-FlowNet (Tian and Andrade-Cetto, 2022) 0.39 1.47 0.57 2.08 1.20 3.99 0.95 3.13
ST-FlowNet1 (ours) 0.40 1.24 0.48 1.86 0.89 2.98 0.70 2.34 1.37 0.92 1.48 0.90
SNN LIF-EV-FlowNet (Hagenaars et al., 2021) 0.53 2.02 0.71 2.63 1.44 4.93 1.16 3.88 1.21 0.95 1.24 0.94
XLIF-EV-FlowNet (Hagenaars et al., 2021) 0.45 1.67 0.73 2.72 1.45 4.93 1.17 3.91 1.23 0.95 1.25 0.93
LIF-FireNet (Hagenaars et al., 2021) 0.57 2.12 0.98 3.72 1.77 6.27 1.50 5.23 1.28 0.99 1.34 1.00
XLIF-FireNet (Hagenaars et al., 2021) 0.54 2.07 0.98 3.73 1.82 6.51 1.54 5.43 1.29 0.99 1.39 0.99
ST-FlowNet2 (ours) 0.37 1.24 0.50 1.86 0.84 2.78 0.70 2.34 1.37 0.91 1.47 0.90
ST-FlowNet3 (ours) 0.39 1.47 0.51 1.92 0.99 3.33 0.77 2.56 1.34 0.93 1.45 0.91
4 Experiments
In this section, we first provide a comprehensive overview of our experimental settings, encompassing datasets, evaluation metrics, and implementation details. Next, we present an in-depth performance comparison between ST-FlowNet and other state-of-the-art models on diverse benchmark datasets. For the purpose of visualization, we display representative examples for qualitative illustration. Finally, we conduct a series of ablation studies to demonstrate the significance of the proposed components.
4.1 Experimental Settings
4.1.1 Optical Flow Datasets
We train the models using the UZH-FPV drone racing dataset (Zhu et al., 2018b), which is distinguished by a diverse distribution of optical flow vectors. We evaluate the models’ performance using the Event-Camera Dataset (ECD) (Mueggler et al., 2017), Multi-Vehicle Stereo Event Camera (MVSEC) (Zhu et al., 2018b, a), and High-Quality Frames (HQF) (Stoffregen et al., 2020) dataset, all captured in real-world scenarios using various DAVIS neuromorphic cameras. Tab. 1 shows a detailed overview of the dataset.
4.1.2 Evaluation Metrics
For the MVSEC (Zhu et al., 2018b) optical flow dataset, the ground truth is generated at each APS frame timestamp and scaled to represent the displacement for the duration of one (dt=1) and four (dt=4) APS frames (Zhu et al., 2018a). Consequently, optical flow is also computed at each APS frame timestamp, using all events within the time window as input for dt=1, or $25\%$ of the window events at a time for dt=4. Both predicted optical flows are evaluated using the average endpoint error (AEE1 for dt=1 and AEE4 for dt=4).
The ECD and HQF datasets do not include ground truth, and thus we employ two image compensation quality metrics to estimate predicted flows. Specifically, the flow warp loss (FWL) assesses the sharpness of the image of warped events compared with the original event partition, and the variance of the contrast of the event images is reported (T. Stoffregen, 2020; Gallego et al., 2019, pp. 12280–12289). Additionally, we also report the ratio of the squared average timestamps (RSAT), indicating the contrast of $\mathcal{L}_{\mathrm{contrast}}$ between predicted optical flow and baseline null vectors (Hagenaars et al., 2021).
4.1.3 Implementation Details
We implement ST-FlowNet using the PyTorch framework and execute on an NVIDIA A100 GPU. The training process consists of 100 epochs for the ANN model, followed by 10 epochs of retraining in the BISNN method, with a batch size of 8. We use the adaptive moment estimation optimizer (Kingma and Ba, 2014) with an initial learning rate of 0.0002, subject to exponential decay. We empirically set the scaling weight for $\mathcal{L}_{\mathrm{smooth}}$ to $\lambda=0.001$, as used in previous work (Zhu et al., 2018a).
In the SNN models converted from ANN models, the firing thresholds are determined using the threshold balance strategy. To simplify the parameter selection process, we focus exclusively on the membrane potential decay factors in the Generator ($\tau_{0}$) and ConvGRU2 ($\tau_{1}$) modules, while setting those in other modules to 0. For consistency and a fair comparison, during the retraining procedure of the BISNN method, both the firing threshold and membrane potential decay factors are initialized with the same values as those used in the A2S method.
4.2 Comparison with State-of-the-art Methods
To validate the effectiveness of our proposed ST-FlowNet, we conduct a comprehensive comparative analysis, assessing its performance compared with other state-of-the-art models from both quantitative and qualitative perspectives (Hagenaars et al., 2021; Scheerlinck et al., 2020; Tian and Andrade-Cetto, 2022). The ST-FlowNet model trained using the ANN, A2S, and BISNN methods are denoted as ST-FlowNet1, ST-FlowNet2, and ST-FlowNet3, respectively.
Figure 3: Visual comparison of ST-FlowNet with other models. Original frame-based images, ground truth of the MVSEC dataset, and the color coding of the optical flow are provided for reference. The AEE1 (black) or FWL (red) results of each predicted optical flow are provided at the upper-left.
4.2.1 Quantitative Evaluation
We conduct a comparative analysis between ST-FlowNet and other state-of-the-art models on the MVSEC dataset across four representative scenarios: outdoor_day1 (denoted as OD1), indoor_flying1 (IF1), indoor_flying2 (IF2) and indoor_flying3 (IF3), considering both the dt=1 and dt=4 conditions. As shown in Tab. 2, ST-FlowNet demonstrates superior AEE performance compared with other ANN and SNN models across most of scenarios. Notably, the optical flow estimation performance can be improved slightly by leveraging the converted SNN model, as particularly evident in the OD1 and IF2 scenarios. For the sake of comparison, on the ECD and HQF dataset, we present a comparative analysis of state-of-the-art models without distinguishing scenarios. As shown in Tab. 2, ST-FlowNet attains the optimal performance, excelling with respect to both the FWL and RSAT metrics.
Table 3: Performance analysis of the key modules in the ST-FlowNet architecture on the MVSEC dataset. The number of Decoder modules is denoted as ’#Ds’, while the ConvGRU1 and ConvGRU2 modules are referred to as ’Aug’ and ’Align’, respectively. The invalid results is highlighted in $\underline{underline}$.
Method #Ds Aug Align OD1 IF1 IF2 IF3
AEE1$\downarrow$ AEE1$\downarrow$ AEE1$\downarrow$ AEE1$\downarrow$
ANN 4 $\times$ $\times$ 0.58 0.60 1.07 0.99
1 $\times$ $\times$ 0.45 0.58 1.02 0.84
1 $\checkmark$ $\times$ 0.44 0.53 0.99 0.81
1 $\checkmark$ $\checkmark$ 0.40 0.48 0.89 0.70
A2S 4 $\times$ $\times$ 0.56 0.63 1.12 0.94
1 $\times$ $\times$ 0.48 0.62 1.03 0.81
1 $\checkmark$ $\times$ 0.53 0.54 0.98 0.80
1 $\checkmark$ $\checkmark$ 0.37 0.50 0.84 0.70
BISNN 4 $\times$ $\times$ 0.67 0.82 1.45 1.22
1 $\times$ $\times$ 0.47 0.64 1.13 0.91
1 $\checkmark$ $\times$ 0.49 0.58 1.05 0.83
1 $\checkmark$ $\checkmark$ 0.39 0.51 0.99 0.77
Table 4: Performance analysis of the key modules in the ST-FlowNet architecture on the ECD and HQF datasets. The number of Decoder modules is denoted as ’Ds’, while the ConvGRU1 and ConvGRU2 modules are referred to as ’Aug’ and ’Align’, respectively.
ECD Method #Ds Aug Align D6 B6 P6 SD
[…]
ANN 4 $\times$ $\times$ 1.29 0.90 1.51 0.92 1.45 0.92 1.40 0.92
1 $\times$ $\times$ 1.45 0.87 1.62 0.92 1.61 0.92 1.61 0.90
1 $\checkmark$ $\times$ 1.42 0.88 1.62 0.91 1.59 0.91 1.73 0.89
1 $\checkmark$ $\checkmark$ 1.46 0.87 1.63 0.92 1.61 0.91 1.77 0.89
A2S 4 $\times$ $\times$ 1.24 0.93 1.44 0.94 1.39 0.94 1.30 0.93
1 $\times$ $\times$ 1.40 0.88 1.60 0.92 1.57 0.92 1.51 0.91
1 $\checkmark$ $\times$ 1.40 0.89 1.61 0.92 1.59 0.91 1.56 0.91
1 $\checkmark$ $\checkmark$ 1.44 0.88 1.63 0.92 1.62 0.91 1.62 0.90
BISNN 4 $\times$ $\times$ 1.31 0.90 1.50 0.93 1.47 0.93 1.39 0.89
1 $\times$ $\times$ 1.44 0.88 1.59 0.93 1.57 0.93 1.74 0.91
1 $\checkmark$ $\times$ 1.44 0.88 1.61 0.91 1.59 0.91 1.83 0.89
1 $\checkmark$ $\checkmark$ 1.45 0.88 1.62 0.92 1.61 0.91 1.66 0.89
HQF Method #Ds Aug Align BO EP HTP RM
[…]
ANN 4 $\times$ $\times$ 1.37 0.91 1.29 0.93 1.49 0.90 1.24 0.94
1 $\times$ $\times$ 1.57 0.90 1.57 0.90 1.79 0.89 1.46 0.91
1 $\checkmark$ $\times$ 1.65 0.90 1.58 0.91 1.81 0.88 1.52 0.91
1 $\checkmark$ $\checkmark$ 1.65 0.90 1.60 0.91 1.82 0.88 1.54 0.90
A2S 4 $\times$ $\times$ 1.30 0.93 1.24 0.95 1.37 0.92 1.20 0.95
1 $\times$ $\times$ 1.46 0.91 1.46 0.91 1.64 0.89 1.35 0.92
1 $\checkmark$ $\times$ 1.58 0.91 1.58 0.91 1.79 0.89 1.48 0.91
1 $\checkmark$ $\checkmark$ 1.60 0.90 1.59 0.91 1.79 0.90 1.49 0.90
BISNN 4 $\times$ $\times$ 1.45 0.89 1.37 0.91 1.55 0.88 1.33 0.92
1 $\times$ $\times$ 1.56 0.91 1.57 0.92 1.67 0.91 1.52 0.92
1 $\checkmark$ $\times$ 1.56 0.91 1.59 0.91 1.71 0.89 1.53 0.91
1 $\checkmark$ $\checkmark$ 1.57 0.90 1.56 0.91 1.77 0.88 1.49 0.91
4.2.2 Qualitative Evaluation
In Fig. 3, we present the visualization results of the predicted optical flows generated by various models. We use EV-FlowNet (ANN model) and LIF-EV-FlowNet (SNN model), which exhibit outstanding performance, for comparison with our ANN and SNN models, respectively. The AEE1 (black) or FWL (red) values are provided at the upper-left of each predicted optical flow image. Overall, ST-FlowNet demonstrates evident superiority in the visual quality of its optical flows compared with competing models. Specifically, in challenging scenarios such as boundary regions with motion blur or sparse features (i.e. D6 and HTP), both EV-FlowNet and LIF-EV-FlowNet yield prediction failures or errors. ST-FlowNet excels in accurately capturing detailed scene information, thereby achieving reliable optical flow estimation. Additionally, qualitative comparisons highlight that the converted SNN model retains the original ANN model’s proficiency in optical flow estimation.
4.3 Ablation Analysis
Our novel ST-FlowNet architecture comprises the following principal components: decoders with a non-pyramid architecture (Decoder), spatio-temporal optical flow augmentation  (ConvGRU1), and optical flow alignment (ConvGRU2). To demonstrate the effectiveness of these components, we establish a baseline architecture by removing two ConvGRU layers and retaining only one decoder in ST-FlowNet. We conduct relative ablation analyses on the baseline architecture by progressively introducing the relevant components.
4.3.1 Analysis of the Number of Decoders
We compare the baseline models with varying numbers of decoders to examine the effectiveness of the decoders. Specifically, we increase the number of decoders from one in the baseline model to four in the comparison models. To conduct a thorough comparison, we select four representative scenarios from the ECD and HQF datasets, each denoted as follows: dynamic_6dof (D6), boxes_6dof (B6), poster_6dof (P6), slider_depth (SD), along with boxes (BO), engineering_posters (EP), high_texture_plants (HTP), and reflective_materials (RM). As shown in Tabs. 3 and 4, the performance decreases across all scenarios when more decoders are employed. This suggests that the use of more decoders is not necessarily associated with improved performance. This observation not only alleviates the pressure to increase the model size but also reinforces the advantage of the SNN model in terms of energy consumption.
4.3.2 Analysis of the Spatio-temporal Augmentation
Building on the baseline model with a single decoder, we integrate ConvGRU1 as the spatio-temporal optical flow augmentation component. Unlike the complete ST-FlowNet, the ConvGRU1 layer receives the upsampled basic predicted optical flow as state information. The detailed results are shown in Tabs. 3 and 4. The AEE results for the MVSEC dataset demonstrate that the ConvGRU augmentation layer enhances optical flow estimation performance in most scenarios, except for OD1. The FWL and RSAT results computed for the ECD and HQF datasets also generally corroborate the improvement, which is more pronounced for the HQF dataset. Furthermore, the ConvGRU layers assist in mitigating the performance loss of converted SNN models. In certain scenarios, SNN models exhibit superior performance compared with ANN models, as observed for the IF2 and IF3 scenarios.
4.3.3 Analysis of the Spatio-temporal Alignment
Expanding on the previous model, we continue to introduce ConvGRU2 as a spatio-temporal alignment module. This module is designed to project the basic and historical predicted optical flow into a standardized state space. By doing so, it not only enhances the precision of optical flow prediction but also establishes an effective reference state for input data. As illustrated in Tabs. 3 and 4, the optical flow estimation performance further improves in most scenarios. Moreover, in comparison to the spatio-temporal augmentation module, the alignment module demonstrates superior effectiveness.
Figure 4: The performance comparison of SNN models initialized with different combinations of membrane potential decay factors. (a) SNN models trained using the A2S method. (b) SNN models trained using the BISNN method. The optimal results are highlighted within red boxes.
Table 5: Performance analysis of SNNs trained with different method.
MVSEC Method OD1 IF1 IF2 IF3
AEE1$\downarrow$ AEE1$\downarrow$ AEE1$\downarrow$ AEE1$\downarrow$
STBP 0.78 1.00 1.78 1.57
A2S 0.37 0.50 0.84 0.70
BISNN 0.39 0.51 0.99 0.77
ECD D6 B6 P6 SD
[…]
STBP 1.18 0.95 1.31 0.96 1.30 0.96 1.46 0.92
A2S 1.44 0.88 1.63 0.92 1.62 0.91 1.62 0.90
BISNN 1.45 0.88 1.62 0.92 1.61 0.91 1.66 0.89
HQF BO EP HTP RM
[…]
STBP 1.34 0.93 1.16 0.97 1.44 0.93 1.17 0.96
A2S 1.60 0.90 1.59 0.91 1.79 0.90 1.49 0.90
BISNN 1.57 0.90 1.56 0.91 1.77 0.88 1.49 0.91
Figure 5: The energy consumption of SNN ST-FlowNet models relative to ANN models.
4.4 Analysis of Training Methods
To thoroughly assess the efficacy of various training approaches, we also conduct experiments involving the direct training of SNNs using the STBP method. The initialization operations of biological parameters are consistent with those in the BISNN method. As presented in Tab. 5, the performance of models trained using the A2S and BISNN methods significantly outperforms those trained with the STBP method. These findings suggest that direct training of SNNs for optical flow estimation remains a challenging task, and further highlight the competitive performance improvements achieved by both the A2S and BISNN methods.
The performance of SNN models converted using the A2S method is influenced by the selection of biological parameters. In our experiments, the spike firing thresholds are determined using the threshold balance strategy, while the membrane potential decay factors are chosen empirically. To further verify the impact of parameter settings on model performance, we perform a search over the parameters $\tau_{0}$ and $\tau_{1}$ within the $[0,0.8]$ range. Experiments are conducted on models trained using both the A2S and BISNN methods under various parameter combinations, and the average results across all scenarios are presented in Fig. 4. The optimal results for each set of experiments are highlighted within red boxes.
As shown in Fig. 4(a), the performance of the models fluctuates significantly with changes in the membrane potential decay factors. On the MVSEC dataset, parameter combinations along the secondary diagonal tend to yield the best results. However, for the ECD and HQF datasets, smaller membrane potential decay factor combinations appear to be more effective. Overall, we are still unable to summarize a general principle for parameter selection from this sufficiently large search space. These findings underscore the challenges associated with selecting optimal biological parameters. Nevertheless, the performance of models trained using the BISNN method demonstrates notable robustness across varying parameter combinations. These results substantiate our claim that BISNN operates as a parameter-free training method, effectively alleviating the complexities of parameter tuning while consistently delivering robust optical flow estimation performance.
4.5 Analysis of Energy Consumption
Finally, we assess the energy consumption for different ST-FlowNet models. In contrast to the ANN ST-FlowNet model that mainly employs multiply-accumulate operations, the SNN ST-FlowNet model predominantly leverages sparse accumulate (AC) operations as computational units, but excluding the ConvGRU1/2, Encoder1 and Decoder1 modules. The relative energy consumption (REC), denoted as $\eta=\frac{\Phi_{\text{SNN}}}{\Phi_{\text{ANN}}}$, is employed to evaluate the energy-saving advantages of SNN ST-FlowNet models. Here, $\Phi$ represents the theoretical energy consumption as calculated in previous work (Yao et al., 2023).
We conduct experiments on ST-FlowNet models trained using the ANN and A2S methods, respectively, and the results are summarized in Fig. 5. As expected, the energy consumption reduction achieved by the SNN ST-FlowNet model is approximately $60\%$ across the entire networks (Fig. 5 (a)). If we focus only on specific modules that exclusively use AC operations in the SNN ST-FlowNet model, a remarkable reduction (more than 300-fold) in energy consumption can be observed (Fig. 5 (b)). These observations clearly demonstrate the advantage of energy consumption in SNN ST-FlowNet models.
5 Discussion and Conclusion
In this study, we present a novel neural network architecture, termed ST-FlowNet, which incorporates enhanced ConvGRU layers, designed to align and augment spatio-temporal information, thereby improving optical flow estimation based on event-driven inputs. Our experimental results validate the effectiveness of the proposed ST-FlowNet, demonstrating its superior performance compared to other state-of-the-art models. Ablation studies further highlight the critical role of the integrated components in our architecture for achieving performance improvements.
To address the challenges in training SNN models for optical flow estimation, we propose two methods: (1) the A2S method, which generates an SNN model from a pre-trained ANN model, and (2) a novel BISNN strategy, aimed at mitigating the complexities associated with the selection of biological parameters. Both methods significantly enhance model performance when compared to the direct training method using STBP. Notably, the BISNN method demonstrates greater robustness across various initialized biological parameters.
The limitations of the ST-FlowNet model are discussed below. First, we observe performance degradation in the converted SNN models under specific conditions (e.g., IF1 in Tab. 2 and HTP in Tab. 4), which may be attributed to theoretical conversion errors inherent in the A2S process (Deng and Gu, 2021). This issue could be addressed through the development of more efficient conversion methods. Second, the spatio-temporal augmentation module exhibits a less robust ability to improve performance, potentially due to modality bias between the input event data and the optical flow state. A well-designed ConvGRU1 module may further improve the performance of the ST-FlowNet model by reducing modality bias, deserving to be examined in our further studies.
Acknowledgments
This work was supported in part by the National Key Research and Development Program of China under grant 2023YFF1204200, in part by the STI 2030–Major Projects under grant 2022ZD0208500, in part by the Sichuan Science and Technology Program under grant 2024NSFJQ0004 and grant DQ202410, and in part by the Natural Science Foundation of Chongqing, China under grant CSTB2024NSCQ-MSX0627, in part by the Science and Technology Research Program of Chongqing Education Commission of China under Grant KJZD-K202401603, and in part by the China Postdoctoral Science Foundation under grant 2024M763876.